Get Help

Online: Request Help
Phone: (707) 826-HELP (4357)
Email: help@humboldt [dot] edu
In Person: Library 120 • Hours
Reset HSU Password
System Status

Services For

System Status Archive

System Issue start Who is affected What happened
Campus Data Network Mar 24 2015 - 10:30 East bleachers Campus Network Infrastructure Upgrade Project - Phase 1. http://www2.humboldt.edu/its/cni-network-core-upgrade-schedule
Campus Data Network Mar 19 2015 - 11:00 North campus houses (Telonicher, Balabanis, Ceramics, Sculpture) Campus Network Infrastructure Upgrade Project - Phase 1. http://www2.humboldt.edu/its/cni-network-core-upgrade-schedule
KBOX Ticketing System Mar 13 2015 - 22:00 Campus community submitting technology help tickets and ITS support staff using KBOX KBOX is way behind in needed updates. We need to upgrade KBOX in order to be supported by DELL KACE. **UPDATE**The backup was successful. Throughout this week updates will be applied. Service will be available (but possibly slow) throughout the updates except for the occasional reboot. During a reboot the service will be down for about 10 minutes. Wait and try again. *UPDATE 2** Some update steps are causing extended outages, the KBOX is currently being updated and will likely be unavailable through this evening, estimated recovery 8-9 PM *Update 3** KBox is running 5.5 and will be upgraded to 6.0 starting 5 p.m. today, Thursday 3-19-2015
Campus Data Network Mar 18 2015 - 12:17 Library, various houses, Van Matre, Gist Hall, Sci A-E, Student Health Ctr., Student Business Serv, Siemens Hall, Parc Bldg Cut-overs on CNI project May have caused the problem, but it is being investigated UPDATE 1 - Network has been updated and services have been restored.
Campus Data Network Mar 18 2015 - 10:00 Science Complex Network Cutovers Campus Network Infrastructure Upgrade Project - Phase 1. http://www2.humboldt.edu/its/cni-network-core-upgrade-schedule
Other Mar 18 2015 - 07:37 Campus Computers ITS is investigating the problem UPDATE1: We have a solution to resolve the issue. Please contact desktop services.
Other Mar 17 2015 - 17:30 Users trying to use Citrix Based Virtual Apps and Desktops Infrastructure changes are preformed on the Storefront server that presents users with a list of available apps and desktops. UPDATE: System were back up by 7:30 but the job was not complete until 8:20. Users that use the built-in receiver may need to either close the receiver and reopen it or reboot their computer.
Campus Data Network Mar 16 2015 - 12:05 On campus: Anyone attempting to access web pages or services off campus.Off campus: Anyone attempting to access campus services. We are investigating the problem at this time. No work with the CNI project has started. We are establishing contact with our data network vendor. **UPDATE(1)** There was a problem with the CENIC network between the Oakland and Sunnyvale backbone routers. **UPDATE(2)** We are continuing to experience intermittent outages on CENIC. As we learn more we will provide updates. **UPDATE(3)** CENIC is continuing to work with two equipment vendors to restore all services but has mitigated the impact of the outage by moving traffic to other links. Maintaining degraded status until we receive word that all is back to normal **Update(4)** CENIC is continuing to work on this issue. They performed emergency maintenance on several campus links last night and have additional maintenance scheduled at 9:00 AM. This could affect campus internet connectivity. We will continue to update as information becomes available. **Update (5)** Closing at this point. CENIC is continuing to troubleshoot their 100 GB link between Sunnyvale and Oakland, but they have moved all traffic off of that link. They will notify us via a maintenance update before attempting to move traffic back
Campus Data Network Mar 17 2015 - 11:30 Gist Hall Network Customers What Happened: Campus Network Infrastructure Upgrade Project - Phase 1. https://www2.humboldt.edu/its/cni-network-core-upgrade-schedule
Campus Data Network Mar 17 2015 - 15:00 SHC Network Customers Campus Network Infrastructure Upgrade Project - Phase 1. https://www2.humboldt.edu/its/cni-network-core-upgrade-schedule
Other Mar 18 2015 - 08:30 People wishing to download Windows applications from the self-service application library We are patching the myapps service to resolve some outstanding issues.
Campus Data Network Mar 16 2015 - 15:00 Library network customers Campus Network Infrastructure Upgrade Project - Phase 1. https://www2.humboldt.edu/its/cni-network-core-upgrade-schedule
Campus Data Network Mar 13 2015 - 06:00 All campus network users Previously when this was connected it caused a network outage affecting campus network users.
Other Mar 15 2015 - 05:00 All Users Scheduled Monthly Maintenance.
Campus Data Network Mar 11 2015 - 13:30 All Network Users We are currently investigating, check back for updates. UPDATE 1 - Most services have been restored. 2 locations are still down - Bret Hart House and Warren House. TNS is working to restore these services and still investigating the cause of the outage. UPDATE 2 - Bret Hart House and Warren House continue to be down and TNS continues to work to restore these services and investigate the cause of the outage. UPDATE 3 - Services at Bret Hart House and Warren House have been restored.
CMS/Peoplesoft Mar 11 2015 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During our regularly schedule maintenance window a fix to the Maxient hold process will be installed.
OBI Mar 8 2015 - 14:00 OBI Consumers Upgrade from 11.1.1.5.5 to 11.1.1.7 will update status when complete. Update is complete.
Other Mar 4 2015 - 15:42 vLab System Administrators are investigating the problem UPDATE: We are still working on the problem. UPDATE2: See Workarounds UPDATE3 See Workarounds
Other Feb 26 2015 - 11:15 Citrix vlab vapp xen Attempting to bring Fischer to aggregated link
Network Folders Feb 24 2015 - 19:00 All Folders Users RedHat has issued a Security Update for Samba (SMB) services, that will require the File Sharing service on Folders to be restarted. This should merely take a second, however, if you are copying data during the service restart, you might get an error message.
Other Feb 24 2015 - 17:00 Users of ere-ssh A critical security patch will be applied.
Other Feb 21 2015 - 01:00 HSE 93 - Brero, HSE 75 - Sculpture, HSE 74 - Ceramics, HSE 55 - Balabanis, HSE 91 - Hagopian, HSE 71 - Little Apartments , HSE 54 - Telonicher TNS is rebooting the switch
Other Feb 20 2015 - 14:40 All Users Power Related Event on Server Farm Switch
Other Feb 20 2015 - 13:32 Users attempting to run DARS Audits. ** DARS Audits are once again available and all PeopleSoft applications appear to be behaving normally. Please contact the Help Desk if you are experiencing any problems with these applications. ** We are currently investigating, check back for updates.
Campus Data Network Feb 20 2015 - 11:30 Campus Network users Preparation work for the upcoming Common Network Initiative (CNI) project.
Campus Data Network Feb 19 2015 - 16:15 Campus Network users CSU CNI ALU Project requires singlemode fiber. We are converting core singlemode fiber links over to multimode fiber where possible to prepare for the refresh.
CMS/Peoplesoft Feb 18 2015 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During our regularly scheduled maintenance window MP 15.02 will be installed.
Account Center Feb 16 2015 - 13:50 Account Center/Gamma/OBI Datafile corruption occurred on Gamma requiring a full restore of the database. During this process all data on Gamma will be unavailable as well as the web services normally available through Account Center. OBI login capability is effected by this as well. **Restoration still in progress. While you may be able to access Account Center, we are still working to fully restore all services. Further updates will be provided later this morning.** **Commonly used services have been restored for use although it will be until tomorrow morning before all Account Center jobs are running on schedule. If you’re experiencing a disabled account when attempting to login, please contact the Technology Help Desk for assistance. This systat will be updated tomorrow morning when all services are back on schedule.** **All gamma services should be back to normal. If you are experiencing otherwise, please open a desk desk ticket assigned to programmers so we can look into it.**
Other Feb 16 2015 - 11:43 Natural History Museum [Resolved] PGE reports electrical power issues, Techs resolving by approximately 1:15 pm today
Campus Data Network Feb 16 2015 - 11:23 Users in Telonicher TNS is investigating. Resolved, appears to have been a UPS failure.
feedback